OpenStudy (anonymous):

The following limit represents the derivative of some function f at some point (a, f(a)). Select an appropriate f(t) and a lim(t->1) (t^2 +t -2)/(t-1)

ganeshie8 (ganeshie8):

compare it wid the definition : derivative = \(\large \lim \limits_{t \to a}~\dfrac{f(t) - f(a)}{t-a}\)

ganeshie8 (ganeshie8):

can u find some \(a\) and \(f(t)\) that match the given limit ?

OpenStudy (anonymous):

a=-1?

OpenStudy (anonymous):

f(t)=t^2?